OpenStudy (anonymous.):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (shamil98):

point-slope form to slope-intercept form y - y1 = m(x-x1) plug it in then change it to slope-intercept form!

OpenStudy (anonymous):

y - y1 = m (x-x1) y- (-3) = -6 (x - (-2)) y + 3 = -6x -12 y = -6x - 15, A

OpenStudy (anonymous.):

@lilo_stitch and @shamil98 Thank you both I just added it to my notes :3

OpenStudy (anonymous):

urwell